Calculate the distance an arrow will land from a hunter if he releases the arrow
with a speed of 20m/s at a 60 degree angle above the horizontal. The tip of the
arrow is 1.5 m above the ground at the point of release.
1
Expert's answer
2019-07-24T15:48:44-0400

the maximum height at which the arrow will fly relative to its tip


"H=\frac{u^2\times\sin^2(\alpha)} {2\times g}"

u - initial speed;

α - angle to the horizontal;

g -acceleration of gravity (10 m/s^2);

H= 15 m;

the time of flight of an arrow to its maximum height (from tip):


"t=\sqrt{\frac{2 \times H}{g}}"

t1.73 s;

the time of flight of an arrow from the maximum height to the ground:

"T=\sqrt{\frac{2 \times (H+h)}{g}}"

h-the initial height of the tip;

T1.82;

the distance at which the arrow falls:


"L=u\times\cos(\alpha)\times(t+T)"

L35.5m

Answer 35.5m